# Critical behavior and multistability in quasi-two-dimensional turbulence

### v1

@@ Abstract
Figure 3: Nonlinear dynamical model of quasi-2D turbulence. (a) The single-mode model 2 for $N=1$, $d_{l}=$ 0.05, 0.1, and 0.15 and $g=0$, $g>0$ shifts the curve along the y-axis. (b) Stable (full lines) and unstable (dashed lines) stationary solutions of (2) as a function of $g$ (normalized by the position of the local minimum or the inflection point $g_{c}$ of (2)) for the same values of $d_{l}$ as in panel (a). (c) Multistability with several overlapping pairs of saddle-node bifurcations develops by coupling multiple modes of vortex density in a cascading manner as in (2), $N=4$. The orange dashed lines show unstable stationary solutions.
To better understand the intermittent region in device J, we recorded a 32-second-long time series of the response amplitude (an example is shown in appendix Supplementary Information, Fig. S17) for 1.45 K, 1.65 K, and 1.85 K at a set of fixed drive amplitudes. Since the timescale of the forcing is set by the flow oscillation period (about 0.5 ms, see Fig. 1(b)), even relatively short datasets allow the study of the long-time behavior of the flow. The resulting two-dimensional histogram in Fig. 4(a) shows a bimodal shape equivalent to the ramped drive data in Fig. 2. The observed time series was threshold-binarized to detect transitions which were then used to calculate the lifetimes of the more dissipative (lower velocity) state as the average time between the system jumping into and out of the state. The probability density function of the state lifetime is exponential (see appendix Supplementary Information Fig. S16), suggesting that the transitions are memoryless. The mean lifetime $\tau$ as a function of the velocity, shown in Fig. 4(b), shows a diverging behavior, which can be captured by a power law of the form $\tau\propto(v_{c}-v)^{-\eta}$ with $\eta\approx 1.33$ (inset of Fig. 4(b)).
@@
Figure 4: Critical behavior of turbulent transition. (a) Histograms of observed velocity amplitude for a fixed drive at 1.45 K, 1.65 K, and 1.85 K (left to right) for device J. The rug plot indicates the used drives. (b) The mean lifetime of the more dissipative (lower velocity) state as a function of velocity in that state for the three temperatures shifted by the critical velocity (indicated in the histograms by horizontal dashed lines). (inset) Lifetime against reduced dimensionless velocity $(v_{c}-v)/v_{c}$ in logarithmic scale. The black dashed line indicates a power-law with a critical exponent of 1.33.
The appearance of the fluctuating regime in device J resembles reversible-irreversible transitions in periodically driven many-body systems Reichhardt et al. (2023). In particular, the time to reach a steady state in periodically driven systems exhibiting random organization (i.e., a dynamical phase transition into a self-organized state through chaotic dynamics) displays the same critical exponent Corté et al. (2008) as the lifetimes of the turbulent states in Fig. 4. The exponent is also close to the longitudinal correlation critical exponent of 2+1D directed percolation $\xi_{||}\approx 1.295$ Avila et al. (2011); Manneville (2016); Hof (2023). The essential feature of our dynamical model (2) describing the formation of multistability is the decrease in the rate of vortex annihilation at high densities, which is consistent with the spatial separation of vortices of opposite signs, that is, the formation of a stable pattern, possibly driven by the inverse turbulent cascade. Similar clustering was observed in the transient evolution of few-vortex systems in Bose-Einstein condensates Gauthier et al. (2019), superfluid helium Sachkou et al. (2019), and numerically in systems with odd viscosity de Wit et al. (2024), although the dynamics of this clustering in steady-state driven turbulence remain numerically largely unexplored.
Our results show that periodically driven quasi-2D flows are closely related to a broad class of systems displaying reversible-irreversible transitions Reichhardt et al. (2023) despite being turbulent and with a fluctuating number of constituent vortex points. The aspect ratios of the studied flows are comparable to geophysical flows, and thus, the presented results suggest the possibility that similar self-organizing behavior occurs in large-scale planetary flows, such as the hexagonal structure in the polar region of Saturn Yadav and Bloxham (2020), or in the atmosphere of Jupiter Marcus (2004) and will follow similar critical behavior as shown in Fig. 4. Indeed, self-organization due to the inverse cascade was recently shown numerically for a dry atmosphere Alexakis et al. (2024). The observed saddle-node bifurcations give experimental support to the recently raised possibility of the collapse of the Atlantic meridional overturning circulation Ditlevsen and Ditlevsen (2023); van Westen et al. (2024) with significant consequences to the global climate.
@@ Acknowledgements.
Supplementary Information
This supplemental information contains additional details on the characterization of nanofluidic devices, derivation of the relationship between vortex density and fourth sound damping, numerical simulations related to confined superflow, and plots showing the complete dataset whose subset was shown in Fig. 2 in the main text.
Appendix A Nanofluidic Helmhholtz resonators
The flows studied are realized in nanofluidic Helmholtz resonators similar to those used in refs. Rojas and Davis (2015); Souris et al. (2017); Shook et al. (2020); Varga and Davis (2021); Varga et al. (2022). A semi-open nanofluidic channel is etched onto a fused silica substrate, two of which are then bonded together to form a fully enclosed nanofluidic cavity approximately 500 nm in height (see Rojas and Davis (2015); Souris et al. (2017) for fabrication details). The cavity consists of a circular basin (5 mm in diameter) and two channels connecting the basin with the bulk 4He (see the schematic of a resonator in Fig. S1(a)). The characteristic length scale at which a viscous fluid follows the motion of an oscillating wall is the viscous penetration depth $\delta=\sqrt{2\nu/\omega}$, where $\nu$ is the kinematic viscosity and $\omega$ the flow oscillation frequency. For the typical resonance frequency 2 kHz and temperature 1.65 K, and using the normal fluid kinematic viscosity from Donnelly and Barenghi (1998), $\delta\approx$2.7 $\mu$m, which is 5-times greater than the spacing between the upper and lower walls of the cavity. We can, therefore, consider the viscous normal fluid component of superfluid helium to be at rest. Four different chip geometries were used in the study, shown in Fig. S2, which provide different velocity profiles of the superfluid flow inside the channel. The geometry was designed to obtain similar Helmholtz resonance frequencies.
Fig.  S1: Nanofluidic resonator and its response. (a) Sketch of the Helmholtz resonator with typical dimensions. The transparent gray represents the cavity filled with superfluid 4He. The aluminum electrodes cover the entire top and bottom wall of the nanofluidic cavity with a 10 µm gap between the side walls and the edge of the electrode. The substrate enclosing the nanofluidic volume is not shown. (b) A photograph of an assembled Helmholtz resonator used for measurement next to a 5 cent Euro coin for scale. (c) The capacitance bridge circuit used for the drive and detection of the flow. The center-tapped transformer excites both sides of the bridge circuit with an alternating voltage with $\pi$ phase shift between the bridge arms. When the bridge is tuned, only the current resulting from the oscillation of the device capacitance flows through the lock-in amplifier. In order to avoid measuring at the second harmonic (since the electrostatic force is proportional to voltage squared), the device is biased by (typically) a 10 V DC voltage. (d) Total detected current magnitude against frequency for several driving voltage amplitudes. For a sufficiently strong drive, the peak becomes non-Lorentzian with a flattened top.
@@ Appendix A Nanofluidic Helmhholtz resonators
The flow in the resonator is driven by two aluminum electrodes deposited opposite each other inside the cavity. An applied alternating voltage causes deformation of the substrate, leading to pressure oscillation inside the cavity and, thus, to a pressure gradient along the channels. The resulting motion of the superfluid causes a fluctuation in the capacitance of the device $\delta C(t)$. This fluctuating capacitance is detected using a bridge circuit shown in Fig. S1(c). The capacitance of the lower branch $C$ is adjusted to the static capacitance of the device $C_{0}$ (typically about 400 pF). Sufficiently far from the Helmholtz resonance, zero current is measured by the lock-in amplifier when the bridge is tuned. As the capacitance fluctuations increase close to the resonance, the charge on the device starts to oscillate due to the bias, which is detected by the lock-in amplifier as an oscillating current.
Fig.  S2: Geometry of the resonator channels. The height of all four cavities is 500 nm. The aluminum electrodes are not shown.
The frequency $f_{0}$ of the Helmholtz resonance is given by Souris et al. (2017)
$f_{0}=\dfrac{1}{2\pi}\sqrt{c\dfrac{\rho_{s}ak_{p}}{l\rho^{2}A^{2}(1+\Sigma)}},$ (S1)
@@
where $\Sigma=(\chi Dk_{p})/(2A)$, $A$ is the area of the basin, $a$ the cross-sectional area of the channel and $l$ its length, and $\chi$, $\rho$ and $\rho_{s}$ are isothermal compressibility, and total and superfluid densities. Compresibility values $\chi$ are taken from Brooks and Donnelly (1977), the total density $\rho$ and the superfluid density $\rho_{s}$ from Donnelly and Barenghi (1998). The parameter $c$ is one of the geometric correction parameters introduced by Souris et al. Souris et al. (2017) that corrects for the finite flow in the basin. The measured resonance frequencies for the four devices, along with a fit to (S1), are shown in Fig. S3. Furthermore, we find an improvement of the fit if we allow for a small systematic error in the temperature calibration as $T_{r}=bT$, where $T$ is the measured temperature, $T_{r}$ is the actual temperature, and $b$ is the correction parameter. The correction parameters are shown in Tab. 1.
@@
Fig.  S3: (a) Resonance frequencies for temperatures between 1.325 K and 1.950 K for the four used chip geometries. The dashed line is fit to (S1) considering the temperature correction; for the fit parameters, see Tab. 1. (b) The corresponding widths (FWHM) of the resonances. The width is equal to the linear part of the damping of the resonance motion. The black dashed line is a fit of the sum of (S26) and (S27).
Chip Type $b$ $c$ $\lambda$
Corner 0.97 1.77 1.94 $\pm$ 0.08
Rounded corner 0.97 2.54 2.05 $\pm$ 0.07
Jet 0.97 1.82 2.07 $\pm$ 0.10
Grid 0.97 2.78 1.77 $\pm$ 0.09
Table 1: Values of correction parameters obtained by fitting Eq. (S1) in case of $b$ and $c$, and the sum of Eq. (S26) and (S27) in the case of $\lambda$. The error of $b$ correction factor is $\pm$ 10-3 and of $c$ correction factor $\pm$ 2$\times$10-3.
Appendix B Surface roughness and the first critical velocity
Due to the stringent requirements of the surface cleanliness for the bonding of the resonators, any manipulation of the unbonded chips outside of the cleanroom environment is essentially destructive, therefore we were unable to measure the topography on the actual devices used in the study. However, a scan of the aluminum layer surface topography using atomic force microscopy is shown in Fig. S4 on devices created using a nearly identical fabrication process. The scan shows 10 µm square on the edge of the aluminum. The step height is 48.64 $\pm$ 0.04 nm, and the RMS surface roughness (that is, the root-mean-square distance to the mean plane) of the aluminum layer is approximately 4.1 nm.
Due to the small size of the vortex core, $a_{0}\approx 0.15$ nm, vortices will be pinned on surface roughness and will be unable to slide unless subject to superflow faster than the depinning velocity, which was estimated by Schwarz Schwarz (1985) for a semi-spherical protuberance of radius $b$ on a planar surface as
$v_{\mathrm{pin}}=\frac{\kappa}{2\pi D}\ln\left(\frac{b}{a_{0}}\right).$ (S2)
Replacing the protuberance size $b$ with the measured surface roughness results for $D=500$ nm in $v_{\mathrm{pin}}\approx$ 11 cm/s, which is within about factor 2 of the initial reversible transitions to nonlinear dissipation seen in Fig. S10 and Figs. S11S14. The origin of the slight temperature dependence of this critical velocity is unknown. However, (S2) was obtained for steady flow; in the present case, the temperature dependence is probably the result of the interaction of the temperature-dependent flow oscillation frequency and the vortex relaxation time given by the mutual friction.
Fig.  S4: Atomic force microscopy scan of the topography of the edge of the aluminum surface (higher, rougher area) in the basin area.
Appendix C Helmholtz resonance and vortex density
Similarly to the second sound Varga et al. (2019), the fourth sound used in this work is damped by the presence of quantized vortices, and its damping can be used to calculate the average number of vortices per unit area (vortex density $L$). For completeness, we briefly reproduce the derivation of the equation of motion of the Helmholtz mode and the damping due to quantized vortices based on refs. Varga et al. (2020); Novotný et al. (2024).
$\delta\rho=\delta\left(\dfrac{M_{B}}{V_{B}}\right)=\dfrac{\delta M_{B}}{V_{B}}-\dfrac{M_{B}}{V_{B}^{2}}\delta V_{B}=\dfrac{1}{V_{B}}(-2a\rho_{s}y+2\rho Ax),$ (S3)
$\delta P=\dfrac{1}{\rho\chi V_{B}}(2\rho Ax-2a\rho_{s}y).$ (S4)
$F=k_{p}x+A\delta P,$ (S5)
where $k_{p}$ is the spring constant and we neglect the inertia of the basin walls. Combining (S3),(S4),(S5), we write the equation of motion for the fluid displacement $y$ in the channel with length $l$
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}}{\rho}a\delta P-F_{f},$ (S6)
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}ak_{p}}{\rho^{2}(\chi V_{B}k_{p}+2A^{2})}\left[\dfrac{2A\rho}{k_{p}}F-2a\rho_{s}y\right]-F_{f}.$ (S7)
Here $F=C_{0}U_{0}(t)U_{B}/D$, since the total voltage is a sum of the bias voltage $U_{B}$ and the driving alternating voltage $U_{0}(t)$ and we only take the component that is resonant with the Helmholtz mode. The friction force $F_{f}$ can be divided into two contributions $F_{f}=F_{\mathrm{in}}+F_{\mathrm{ns}}$. The first part is the intrinsic damping of the superfluid motion, which is discussed in Section E, and can be written in the form
$F_{\mathrm{in}}=al\rho_{s}\gamma_{0}\dot{y},$ (S8)
where $\gamma_{0}$ is a dimensionless intrinsic damping factor. The second contribution is the mutual friction force between the superfluid and quantized vortices, which can be expressed Donnelly (1991) as
$\mathbf{F}_{\mathrm{ns}}=al\alpha\rho_{s}\hat{\bm{\omega}}\times[\bm{\omega}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}})]+al\alpha^{\prime}\rho_{s}\bm{\omega}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}}),$ (S9)
where $\mathbf{v_{\mathrm{n}}}$ and $\mathbf{v_{\mathrm{s}}}$ are velocities of the normal and the superfluid component. Macroscopic vorticity $\bm{\omega}$ represents the mean circulation of a set of parallel-oriented vortices, and $\alpha,\alpha^{\prime}$ are the temperature-dependent mutual friction coefficients Donnelly and Barenghi (1998). Due to geometrical constraints we have for the vortex density $L\approx\omega/\kappa$, where $\kappa\approx 9.97\times$10-8 m2s-1 is the quantum of circulation.
Assuming that the normal fluid component is viscously clamped to the walls, i.e. $\mathbf{v_{\mathrm{n}}}=0$, and neglecting the second non-dissipative mutual friction term, (S9) simplifies to
$F_{ns}=al\alpha\kappa\rho_{s}Lv_{\mathrm{s}}\sin^{2}{\theta},$ (S10)
where the angle $\theta$ is between $\mathbf{v_{\mathrm{s}}}$ and $\hat{\bm{\omega}}$. Assuming that the vortices are preferentially oriented perpendicularly to the superfluid flow (a consequence of 2-dimensionality), we have $\sin^{2}{\theta}=1$.
The dynamics of $L(t)$ likely occurs on a significantly longer time scale than the flow oscillation period, as evidenced by the relatively long-lived states in both the multi-stable and intermittently switching regimes (from roughly the inverse linewidth up to several seconds). In the steady state, we therefore consider $L$ to be a constant. Note, however, that this constant will depend on the RMS flow velocity amplitude (and thus drive amplitude and frequency); therefore, a Lorentzian peak shape should not be expected. A similar constancy of $L$ generated by the second sound in a 3D geometry was reported in Midlik et al. (2021).
Identifying $\dot{y}$ with $v_{\mathrm{s}}$ and substituting friction (S8),(S10) into the equation of motion (S7) results in
$\ddot{y}+(\alpha\kappa L+\gamma_{0})\dot{y}+\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})}y=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t),$ (S11)
$\ddot{y}+\tilde{\gamma}\dot{y}+\omega^{2}_{0}y=f_{0}e^{i\omega t},$ (S12)
$\tilde{\gamma}$ $=\alpha\kappa L+\gamma_{0}=\gamma(L)+\gamma_{0},$ (S13)
$\omega_{0}^{2}$ $=\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})},$ (S14)
$f_{0}(t)$ $=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t)=f_{0}e^{i\omega t}.$ (S15)
$\dfrac{y_{0}}{y}-1=\dfrac{\gamma(L)}{\gamma_{0}}.$ (S16)
Using Eq. S10 (for $\theta=\pi/2$) and the relation between the damping factor $\gamma_{0}$ and the width $\Delta_{0}$ of the un-damped resonance curve, $\gamma_{0}=2\pi\Delta_{0}$, results in
$L=\dfrac{2\pi\Delta_{0}}{\alpha\kappa}\left(\dfrac{Y_{0}}{Y}-1\right).$ (S17)
In the present work, $y_{0}$ is extrapolated from the linear regime of the force-velocity curves shown in Fig. 1D of the main text and Fig. S10.
Appendix D Calibration of force and velocity

### v2

@@ III.2 Hysteresis and multistability in turbulent transition
Figure 5: Nonlinear dynamical model of quasi-2D turbulence: (a) The single-mode model in Eq. (18) for $N=1$, $d_{l}=$ 0.05, 0.1, and 0.15 (top to bottom curve) and $g=0$, $g>0$ shifts the curve along the y-axis. (b) Stable (full lines) and unstable (dashed lines) stationary solutions of (18) as a function of $g$ (normalized by the position of the local minimum or the inflection point $g_{c}$ of (18)) for the same values of $d_{l}$ as in panel (a). (c) Multistability with several overlapping pairs of saddle-node bifurcations develops by coupling multiple modes of vortex density in a cascading manner as in (18), $N=4$. The orange dashed lines show unstable stationary solutions.
@@
Figure 6: Critical behavior of turbulent transition in device J: (a) Time dependence of the observed velocity amplitude in the intermittently switching region at 1.65 K at drive pressure gradient approximately 9.5 kPa/m. Maximum and minimum velocity were normalized to 0 and 1, respectively; the orange line shows the binarized signal used for the detection of transitions. (b) Histrogram of the turbulent state lifetimes in the transitional region at the three indicated temperatures for datasets with a mean lifetime in the range 0.1 to 0.2 s. A straight line in semilogarithmic axes indicates exponentially decaying dependence. (c) Histograms of observed velocity amplitude for a fixed drive at 1.45 K, 1.65 K, and 1.85 K (left to right). The rug plot indicates the used drives. (d) The mean lifetime of the lower (more dissipative) state as a function of dimensionless reduced velocity in that state for the three temperatures (the critical velocity is indicated in the histograms by horizontal dashed lines). The power law (dashed lines) was fit to the range indicated by the shaded areas. (inset) Same data against dimensional shifted velocity $v-v_{c}$ (in m/s).
III.3 Stochastic transitions and critical behavior.
In the region of intermittent transitions in device J it is apparent that the the probability of finding the turbulence in the more dissipative state gradually increases as the velocity is increased. To capture the statistics of the intermittent transitions for device J we recorded the time series of the response amplitude for 1.45 K, 1.65 K, and 1.85 K at a set of fixed drive amplitudes. An example is shown in Fig. 6(a), which shows clear switching between distinct states; records were 32 s long, i.e., capturing over 60000 flow oscillation periods, which are about 0.5 ms long (i.e., 2 kHz resonance frequency, Fig. 2(b)). The resulting two-dimensional histogram in Fig. 6(c) shows a bimodal shape equivalent to the ramped drive data in Fig. 4. The observed time series was threshold-binarized to detect transitions which were then used to calculate the lifetimes of the more dissipative (lower velocity) state as the average time between the system jumping into and out of the state. The probability density function of the state lifetime is exponential (see Fig. 6(b)), suggesting that the transitions are memoryless. The mean lifetime $\tau$ as a function of the velocity, Fig. 6(d), shows a diverging behavior consistent with a power law of the form $\tau\propto(v_{c}-v)^{-\eta}$ with $\eta=1.35\pm 0.07$. The exponent was calculated as the mean of the three temperatures shown in Fig. 6(b) weighted by fit error determined by bootstrap [46]. We note, however, that the scaling is observed only for approximately half a decade for the two higher temperatures due to a limited number of transitions very close to the critical point. For 1.45 K, we observe only a possible indication of the scaling restricted to $(v_{c}-v)/v_{c}>0.1$; close to the critical point the transitions become less distinguishable.
IV Discussion and conclusions
The critical exponent of the stochastic fluctuating transition is close to the longitudinal correlation critical exponent of 2+1D directed percolation $\xi_{||}\approx 1.295$ which describes transition to turbulence in shear flows [47, 48, 19] although the present layer-height critical $\mathrm{Re}\approx 5$ is significantly smaller than that of classical shear flows $\mathrm{Re}_{c}\approx 330$ [40]. The mechanism for transition to turbulence is, therefore, most likely different between the oscillating superflow and classical shear flows.
The essential feature of our dynamical model (18) describing the formation of multistability is the decrease in the rate of vortex annihilation at high densities, which is consistent with clustering of vortices of equal signs which occurs during vorticity condensation [49, 50]. The intermittent transitions seen in Fig. 6(a) resemble the stochastic transitions between large-scale vortices and zonal jets [7, 5] or between multiple zonal jet configurations [51] observed in several numerical works (including three-state multistability observed in a simulation of anisotropic 2D turbulence by Xu et al. [7]). However, Xu et al. observed that the lifetimes of both types of condensates increase exponentially with increasing Reynolds number, whereas we observe a diverging increase in the lifetime of one state at the expense of the other with a critical exponent. Interestingly, the reversible-irreversible transition in periodically driven many-body systems [52] shows a similar critical exponent, about 1.33 [18], in the time to reach a steady state close to the transition. This suggests that the switch in the preferred turbulent state as flow parameters are varied falls within the universality class of this nonequilibrium phase transition, although measurement of the state lifetimes over a broader range of reduced velocities is needed for a more reliable estimate of the critical exponent. Superficially the
Relaxing the assumption of strict two-dimensionality, the condensate was found to form discontinuously at a critical forcing scale to layer thickness ratio [53, 9, 54, 42]. The geometry of our flow is fixed; however, the typical length scale over which vortices move will rapidly increase upon exceeding the depinning velocity (17), which will result in the increase of the energy injection scale. Indeed, the bistability we observe is similar to scenario 3 of ref. [9], where the formation of 2D inverse cascade and cessation of 3D direct cascade occur at distinct critical points with a bidirectional cascade in the intermediate range, although no power-law divergence was observed there. We are, therefore, unable to definitively identify the nature of the multiple turbulent states and the transitions between them since none of the currently available numerical results fully account for the observed multistability and the statistics of transitions.
Finally, we note in passing that the aspect ratios of the studied flows are comparable to geophysical flows, where self-organizing behavior is common [55, 56, 57] and thus, the presented results might help shed light on transitions and formation of large-scale structures in planetary flows. Indeed, baroclinic turbulence was shown to be well described by a gas of point vortices [58, 59] and multistability is present in the climate model [60], e.g. in the simulations of large-scale flows of the Atlantic ocean [61, 62] in context of the collapse of the Atlantic meridional overturning circulation, which would have significant impact on the global climate.
Data Availability
All data and software needed to generate the figures in the paper are available in Zenodo repository at doi.org/10.5281/zenodo.14824085, which also contains tools to view all underlying data of the histograms shown in Figs. 4 and 6.
@@ Acknowledgements.
Appendix A Drive and detection circuitry
@@
The flow in the resonator is driven by two aluminum electrodes deposited opposite each other inside the cavity. An applied alternating voltage causes deformation of the substrate, leading to pressure oscillation inside the cavity and to a pressure gradient along the channels. The resulting motion of the superfluid causes a fluctuation in the capacitance of the device $\delta C(t)$. This fluctuating capacitance is detected using a bridge circuit shown in Fig. 7(b). The capacitance of the lower branch $C$ is adjusted to the static capacitance of the device $C_{0}$ (typically about 400 pF). Sufficiently far from the Helmholtz resonance, zero current is measured by the lock-in amplifier when the bridge is tuned. As the capacitance fluctuations increase close to the resonance, the charge on the device starts to oscillate due to the bias, which is detected by the lock-in amplifier as an oscillating current.
Figure 7: Nanofluidic resonator and detection circuit. (a) A photograph of an assembled Helmholtz resonator used for measurement next to a 5 cent Euro coin for scale. (b) The capacitance bridge circuit used for the drive and detection of the flow. The center-tapped transformer excites both sides of the bridge circuit with an alternating voltage with $\pi$ phase shift between the bridge arms. When the bridge is tuned, only the current resulting from the oscillation of the device capacitance flows through the lock-in amplifier. In order to avoid measuring at the second harmonic (since the electrostatic force is proportional to voltage squared), the device is biased by (typically) a 10 V DC voltage.
The frequency $f_{0}$ of the Helmholtz resonance is given by [24]
$f_{0}=\dfrac{1}{2\pi}\sqrt{c\dfrac{\rho_{s}ak_{p}}{l\rho^{2}A^{2}(1+\Sigma)}},$ (19)
@@
where $\Sigma=(\chi Dk_{p})/(2A)$, $A$ is the area of the basin, $a$ the cross-sectional area of the channel and $l$ its length, and $\chi$, $\rho$ and $\rho_{s}$ are isothermal compressibility, and total and superfluid densities. Compresibility values $\chi$ are taken from [63], the total density $\rho$ and the superfluid density $\rho_{s}$ from [25]. The parameter $c$ is one of the geometric correction parameters introduced by Souris et al. [24] that corrects for the finite flow in the basin. The measured resonance frequencies for the four devices, along with a fit to (19), are shown in Fig. 8. Furthermore, we find an improvement of the fit if we allow for a small systematic error in the temperature calibration as $T_{r}=bT$, where $T$ is the measured temperature, $T_{r}$ is the actual temperature, and $b$ is the correction parameter. The correction parameters are shown in Tab. 1.
@@
Figure 8: (a) Resonance frequencies for temperatures between 1.325 K and 1.950 K for the four used chip geometries. The dashed line is fit to (19) considering the temperature correction; for the fit parameters, see Tab. 1. (b) The corresponding widths (FWHM) of the resonances. The width is equal to the linear part of the damping of the resonance motion. The black dashed line is a fit of the sum of (28) and (29).
[…]
Corner 0.97 1.77 1.94 $\pm$ 0.08 1.24 338.6
Rounded corner 0.97 2.54 2.05 $\pm$ 0.07 1.19 312.9
Jet 0.97 1.82 2.07 $\pm$ 0.10 1.45 335.7
Grid 0.97 2.78 1.77 $\pm$ 0.09 1.62 466.4
[…]
Appendix B Calibration of force and velocity